OpenStudy (anonymous):

measure of angle A equals 8x minus 2,measure of angle B equals 2x minus 8, andmeasure of angle C equals 94 minus 4x. List the sides of triangle ABC in order from shortest to longest. side AC, side AB, side BC side Ac, side BC, side AB side AB, side AC, side BC side AC, side BC, side AB

OpenStudy (anonymous):

The sum of all the angles will be 180 as it always is in any triangle, so: (8x - 2) + (2x - 8) + (94 - 4x) = 180 Can you solve this for "x"?

OpenStudy (anonymous):

not sure how to sole for x i knoe i need to get x by its self but not sure how to do it

OpenStudy (anonymous):

I'll help you through it.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

(8x - 2) + (2x - 8) + (94 - 4x) = 180 -> (8x + 2x - 4x) + (94 - 2 - 8) = 180 -> 6x + 84 = 180 -> 6x = 96 -> x = 16 So, A = 8(16) - 2 = 126 B = 2(16) - 8 = 24 C = 94 - 4(16) = 30 Now, the shortest side will be across from the smallest angle. The largest side will be across from the largest angle. I'll draw a picture while you study how I did the problem up to this point.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):



OpenStudy (anonymous):

AC is the smallest side AB is the medium-length side BC is the longest side. And that's all there is to it!

OpenStudy (anonymous):

is the hypotnuse always the longest side

OpenStudy (anonymous):

The hypotenuse is the angle opposite the 90 degree angle in a right triangle. So, in a right triangle, the hypotenuse will be the longest side.

OpenStudy (anonymous):

oh okay thank you i need to study this some more but you helped me see how to set up the problem

OpenStudy (anonymous):

You're welcome!